# Decoding Visual Experience and Mapping Semantics through Whole-Brain Analysis Using fMRI Foundation Models

### v2

@@ 4.2 WAVE Model
In this study, we introduce a novel approach to fMRI decoding by proposing a diffusion prior integrated within a flexible diffusion model during the training phase. We design the diffusion prior with adjustable parameters, ensuring these are the only elements subject to learning while training the diffusion part. Distinctly, the training of the diffusion model progresses independently from the contrastive learning process. This method leverages latent fMRI features $\mathcal{H_{\textit{fMRI}}}$, extracted by an fMRI encoder, which has been refined via contrastive learning across three modalities. To streamline our model, we eliminated the text and image encoders, retaining solely the fMRI encoder that produces latent $\mathcal{H_{\textit{fMRI}}}$ in $\in\mathbb{R}^{768}$. The foundation of our diffusion prior is based on UNet [46], which utilizes the fMRI latent space as a conditional input. The primary function of this prior is to transform the fMRI latent into image latent space suitable for further processing. During the training, we incorporate a singular linear layer with weights $W$ to adjust the dimensions of the fMRI latent, effectively transforming them as $\mathcal{H_{\textit{fMRI}_{\text{T}}}}$ = $W\times\mathcal{H_{\textit{fMRI}}}$, with the transformed latent space having dimensions $\mathcal{H_{\textit{fMRI}_{\text{T}}}}\in\mathbb{R}^{257\times 768}$. Our training process employs an $\ell_{2}$ loss $\mathcal{L_{D}}$ to measure the discrepancy between the ground-truth image latent $\mathcal{H_{\textit{image}}}$, encoded by a pre-trained CLIP image encoder, and the image latent $\mathcal{H_{\textit{image}_{\text{D}}}}$ denoised by our diffusion prior from the fMRI latent $\mathcal{H_{\textit{fMRI}_{\text{T}}}}$. We enhance the robustness of our model by applying intensive data augmentation techniques to the input fMRI data, including sequence permutation, normalization, and sparse coding. Furthermore, as a regularization strategy, we adjust the impact of the loss by applying a loss multiplier $\alpha$ of 0.3 as $\alpha\mathcal{L_{D}}$, optimizing the model’s performance and generalization capability.
@@
Following previous works [10, 16] measurement for reconstructed images, we used the 50-way-top-1 accuracy classification task to evaluate the semantic correctness of the reconstructed images. As the same process as Chen et al., we used a pre-trained  [13, 38] ImageNet1K classifier. Both ground-truth and reconstructed images were input to the classifier, and we checked if the top-1 classification of the generated image matched the ground-truth image classification. This metric straightforwardly examines the reconstructed image’s semantic meaning through classification accuracy. This metric is the most critical since the model is aiming to recover the semantic meaning of the images in this work.
In addition to evaluating reconstructed image semantic meaning, we followed previous works by adopting both low-level and high-level metrics as supplementary evaluation for reconstructed image detail features. In low-level metrics, we used pixel-level correlation(PixCorr) to evaluate the reconstructed images and ground-truth images. We used two-way identification to compare the second and fifth layers of AlexNet, and the last pooling layer of InceptionV3, and the output layer of the CLIP-Vision model. EffNet-B (Eff) and SwAV-ResNet50 (SwAV) are used to calculate the distance correlation. Please see Ozcelik and VanRullen for more details.
@@ 4.3 Whole-Brain Visual Experience Analysis
We propose a self-supervised method for analyzing the semantic mapping of whole-brain responses to distinct groups of stimuli and identifying regions of interest. This method specifically examines the correlation between brain region activities and visual images. For a detailed cluster analysis, the self-supervised approach categorizes these images into five distinct clusters. This clustering enhances our understanding of the nuanced relationships between visual stimuli and brain responses.
Clustering
K-Means algorithm is a widely known algorithm for self-supervised clustering. However, K-Means is unreliable when it comes to high dimensional data, i.e., it performs poorly when used directly on images. Hence, we used CLIP Image encoder to encode the image to the image latent space and implemented K-Means on the image latent data as shown in Fig. 3. We clustered the dataset visual stimulus images to 5 clusters. We visualized the image labels in each cluster in a word cloud, where larger word sizes indicated higher word frequency in the image label. We asked ChatGPT-4 of the 5 clusters’ word clouds, and let ChatGPT summarize the word cloud cluster name. ChatGPT summarized the cluster name to: “Pets”, “Outdoor Activities”, “Wild Animals”, “Interiors”, and “Indoor Objects.”
@@
To assess the model’s focus on specific neural networks, we utilized Grad-CAM [50], a gradient-based technique. This method involves tracking the gradients from the target output back through the network layers to the originating fMRI input data. This process helps in pinpointing which areas of the input are most influential in the model’s predictions, where the gradient values are high. In our cluster analysis, the target outputs are the clusters. We created gradient-based saliency map for each cluster. As shown in Fig. 3, we averaged the gradient values from four subjects in the dataset, and we took the top 20 regions for visualization.
@@
During the training and reconstruction phases, we incorporated a visual region mask into the fMRI input data. This mask, aligned with the DiFuMO atlas [11], sets the values within the visual regions to zero by following the network mapping scheme to the Yeo 7 networks [63]. Specifically, we applied the mask to all input fMRI data associated with the visual network as defined in the Yeo 7 classification. This masking procedure was consistently maintained throughout the training and image reconstruction processes. The images reconstructed from the fMRI data with the visual networks masked are displayed in Fig. 2. The ablation network ablation of other six networks are the same process described as visual network above.
@@ 4.4 Zero-Shot Human Imagination Decoding
To assess the quality of the reconstructed images against the ground truth texts, we utilized the Two-Way-Identification Accuracy metric. This involved using a pre-trained CLIP model to encode $n$ reconstructed images and $n$ ground-truth texts into image embeddings $\mathcal{H_{\textit{image}}}$ and text embeddings $\mathcal{H_{\textit{text}}}$, respectively. We computed the Pearson product-moment correlation coefficients [21], resulting in an image-text coefficient matrix $M\in\mathbb{R}^{n\times n}$. The accuracy for each image was determined by the ratio of row elements with values less than their diagonal counterparts, divided by $n-1$. The overall two-way identification accuracy is the average of these individual accuracies.
We recruited 24 young subjects, each presented with 20 scenarios accompanied by text descriptions. Utilizing the aforementioned metric, we aggregated the data into a comprehensive coefficient matrix $M_{\textit{all}}\in\mathbb{R}^{480\times 480}$, and performed a permutation test 5,000 times by randomizing the sequence of image embeddings.
@@
5 Acknowledgement
This research is supported by the HAI-Google Research Award, UST, HAI Hoffman-Yee Grant, and National Institutes of Health grants NIH U24AG072701 and R01AG089169. We also greatly thank Yurong Liu from New York University, for providing figure suggestions and proofreading.
6 Code availability
@@ A.1 Individual Decoding Result
In our main paper, we present two experiment settings, the individual model for decoding and the universal model. To explore our WAVE model decoding ability for individual subjects. We provide each subject metric results here. As shown in the Table 3. We showed all four subjects in the BOLD5000 dataset, and each subject ID is listed in the first column, and the details metrics results are followed in each row. WAVE has great advances in decoding semantic meaning from the fMRI and the results in High-Level and Semantic Top-1 Accuracy also prove this point.
@@
Subject Model Low-Level High-Level Semantic PixCorr↑ SSIM↑ AlexNet(2)↑ AlexNet(5)↑ Incep↑ CLIP↑ Eff↓ SwAV↓ Top-1 Acc↑ CSI1 WAVE (Ours) 0.078 0.230 76.29% 83.00% 75.20% 86.37% 0.859 0.539 33.28% MindEye [49] 0.079 0.226 76.90% 84.06% 67.67% 67.67% 0.895 0.566 17.78% Mind-Vis [10] 0.095 0.313 73.81% 82.18% 74.02% 70.07% 0.877 0.514 23.64% CSI2 WAVE (Ours) 0.061 0.167 72.38% 82.74% 72.06% 82.71% 0.890 0.548 28.53% MindEye [49] 0.046 0.218 67.14% 72.18% 60.99% 68.09% 0.940 0.616 11.01% Mind-Vis [10] 0.076 0.320 69.90% 77.09% 68.39% 70.07% 0.906 0.546 16.04% CSI3 WAVE (Ours) 0.063 0.208 70.14% 78.67% 65.40% 80.04% 0.898 0.550 23.00% MindEye [49] 0.031 0.218 69.95% 73.32% 64.68% 67.53% 0.934 0.611 14.32% Mind-Vis [10] 0.085 0.316 68.66% 74.98% 66.20% 70.07% 0.912 0.553 16.40% CSI4 WAVE (Ours) 0.045 0.193 67.60% 74.38% 62.38% 77.04% 0.922 0.572 16.15% MindEye [49] 0.068 0.207 70.49% 79.69% 71.99% 76.64% 0.895 0.574 24.27% Mind-Vis [10] 0.043 0.286 57.33% 56.73% 59.08% 58.15% 0.961 0.610 14.55%
Visual Network Mask Decoding.
Besides presenting the decoding results from whole brain wise data (WAVE) and visual cortex data (MindEye & Mind-Vis), we also explored decoding information from brain regions that do not belong to the visual networks. By applying the mask to the input fMRI data, WAVE can still successfully decode the natural image from fMRI across all subjects. In addition to the averaged subject results in the manuscript, we also provide the individual visual networks mask results in Table 4. Our individual visual mask result gives new insights into neuroscience to understand various brain regions’ functions not only at visual perception level, but in the cognition. Compared with the whole brain individual decoding results in Fig. 3, the visual masked metrics across subjects are slightly lower.
@@
Subject Low-Level High-Level Semantic PixCorr↑ SSIM↑ AlexNet(2)↑ AlexNet(5)↑ Incep↑ CLIP↑ Eff↓ SwAV↓ Top-1 Acc↑ CSI1 0.059 0.191 67.75% 78.52% 70.12% 82.35% 0.881 0.558 25.95% CSI2 0.053 0.185 67.31% 75.43% 67.42% 78.58% 0.911 0.591 17.40% CSI3 0.033 0.176 65.41% 75.43% 62.49% 75.73% 0.920 0.600 18.78% CSI4 0.041 0.193 66.10% 71.11% 62.62% 75.63% 0.925 0.603 13.79%
@@
We conducted ablation studies to evaluate the performance of different encoder and decoder configurations in diffusion-based image reconstruction. Two types of encoders were used: the pre-trained fMRI Foundation model and the train-from-scratch MLP Backbone. For the decoders, we explored two diffusion models: Versatile Diffusion and Stable Diffusion. The ablation results, presented in Table 5, indicate that the fMRI Model, which employs a region encoder [55], demonstrates superior performance compared to the MLP Backbone. Additionally, Versatile Diffusion consistently outperformed Stable Diffusion across evaluated metrics.
@@
fMRI Encoder Decoder Acc(%) Min(%) Max(%) FID Min Max Region Enc Voxel Enc fMRI Model (Thomas et al.) w/o VD 22.13 ± 1.48 19.62 23.41 2.45 ± 0.71 1.68 3.55 fMRI Model (Thomas et al.) w/o SD 18.81 ± 1.17 17.87 20.46 3.70 ± 1.48 2.37 5.77 MLP Backbone w/o VD 18.79 ± 1.72 16.71 20.92 4.81 ± 0.95 3.53 5.80 MLP Backbone w/o SD 13.03 ± 0.95 12.17 14.35 6.69 ± 1.55 4.94 8.70 w/o MLP Backbone VD 12.91 ± 1.12 11.41 14.09 3.65 ± 0.63 3.09 4.53 w/o MLP Backbone SD 10.41 ± 1.00 9.01 11.23 6.62 ± 1.36 4.87 8.20
In addition to exploring various encoder and decoder configurations, our research extended to analyzing the impact of decoder loss functions on Subject CSI3. Our findings indicate that utilizing diffusion model loss directly on image pixel values, rather than on embeddings, enhances performance in low-level pixel-related metrics and Fréchet Inception Distance (FID). As demonstrated in Table 6, Mind-Vis [10] has better low-level performance.
@@
Model Loss Decoder Acc(%) Min(%) Max(%) FID Min Max WAVE Embed VD 22.13 ± 1.48 19.62 23.41 2.45 ± 0.71 1.68 3.55 MindEye (Scotti et al.) Embed VD 13.45 ± 0.13 12.89 14.97 2.91 ± 0.39 2.53 3.30 Mind-Vis (Chen et al.) Pixel LDM 14.83 ± 2.25 12.59 19.08 1.64 ± 0.19 1.40 1.98
Additionally, we also conduct research to show multi-modal fMRI-Visual-Language learning with prompt Meta-Net effect in this decoding task. As illustrated in Table 7, incorporating multiple modalities through contrastive learning enhances the model’s ability to encode representations, thereby improving the classification accuracy for Brain (fMRI) and Vision (Visual Stimulus) using the CLIP model. The introduction of prompt-based learning further aids in balancing the training across multiple modalities, enhancing both Brain-Vision and Brain-Language accuracy.

### v3

@@ 4.2 WAVE Model
In this study, we introduce a novel approach to fMRI decoding by proposing a diffusion prior integrated within a flexible diffusion model during the training phase. We design the diffusion prior with adjustable parameters, ensuring these are the only elements subject to learning while training the diffusion part. Distinctly, the training of the diffusion model progresses independently from the contrastive learning process. This method leverages latent fMRI features $\mathcal{H_{\textit{fMRI}}}$, extracted by an fMRI encoder, which has been refined via contrastive learning across three modalities. To streamline our model, we eliminated the text and image encoders, retaining solely the fMRI encoder that produces latent $\mathcal{H_{\textit{fMRI}}}$ in $\in\mathbb{R}^{768}$. The foundation of our diffusion prior is based on UNet Ronneberger et al. [2015], which utilizes the fMRI latent space as a conditional input. The primary function of this prior is to transform the fMRI latent into image latent space suitable for generative reconstruction. During the training, we incorporate a singular linear layer with weights $W$ to adjust the dimensions of the fMRI latent, effectively transforming them as $\mathcal{H_{\textit{fMRI}_{\text{T}}}}$ = $W\times\mathcal{H_{\textit{fMRI}}}$, with the transformed latent space having dimensions $\mathcal{H_{\textit{fMRI}_{\text{T}}}}\in\mathbb{R}^{257\times 768}$. Our training process employs an $\ell_{2}$ loss $\mathcal{L_{D}}$ to measure the discrepancy between the ground-truth image latent $\mathcal{H_{\textit{image}}}$, encoded by a pre-trained CLIP image encoder, and the image latent $\mathcal{H_{\textit{image}_{\text{D}}}}$ denoised by our diffusion prior from the fMRI latent $\mathcal{H_{\textit{fMRI}_{\text{T}}}}$. We enhance the robustness of our model by applying intensive data augmentation techniques to the input fMRI data, including sequence permutation, normalization, and sparse coding. Furthermore, as a regularization strategy, we adjust the impact of the loss by applying a loss multiplier $\alpha$ of 0.3 as $\alpha\mathcal{L_{D}}$, optimizing the model’s performance and generalization capability.
@@
Following previous works [Chen et al., 2023, Gaziv et al., 2022] measurement for reconstructed images, we used the 50-way-top-1 accuracy classification task to evaluate the semantic correctness of the reconstructed images. As the same process as Chen et al. [2023], we used a pre-trained  [Dosovitskiy et al., 2020, Paszke et al., 2017] ImageNet1K classifier. Both ground-truth and reconstructed images were input to the classifier, and we checked if the top-1 classification of the generated image matched the ground-truth image classification. This metric straightforwardly examines the reconstructed image’s semantic meaning through classification accuracy. This metric is the most critical since the model is aiming to recover the semantic meaning of the images in this work.
In addition to evaluating reconstructed image semantic meaning, we followed previous works by adopting both low-level and high-level metrics as supplementary evaluation for reconstructed image detail features. In low-level metrics, we used pixel-level correlation(PixCorr) to evaluate the reconstructed images and ground-truth images. We used two-way identification to compare the second and fifth layers of AlexNet, and the last pooling layer of InceptionV3, and the output layer of the CLIP-Vision model. EffNet-B (Eff) and SwAV-ResNet50 (SwAV) are used to calculate the distance correlation. Please see Ozcelik and VanRullen [2023] for more details.
@@ 4.3 Whole-Brain Visual Experience Analysis
We developed a self-supervised analytical framework to map distinct visual semantic categories to their corresponding whole-brain representations. This method examine specific brain region activities covary with distinct clusters of visual images. By partitioning the stimulus set into five semantically coherent clusters, this analysis enhances our understanding of the nuanced relationships between visual stimuli and brain responses.
Semantic Clustering
To investigate the neural correlates of distinct visual concepts, we employed an unsupervised clustering approach on the stimulus set. Since direct clustering of raw pixel values often captures low-level statistics rather than semantic content, we first projected all visual stimuli into a high-dimensional semantic space using the pre-trained CLIP image encoder. We then applied the K-Means algorithm to these latent embeddings to partition the dataset into five distinct semantic clusters (Fig. 3). To interpret the thematic content of each cluster, we aggregated the object labels associated with the images in each group and visualized them as word clouds, where font size corresponds to label frequency. Finally, to derive cohesive category definitions, we utilized a Large Language Model (GPT-4) to synthesize the dominant themes from the label distributions. The resulting category definitions are “Pets”, “Outdoor Activities”, “Wild Animals”, “Interiors”, and “Indoor Objects.” We also manually verified to ensuring they accurately reflected the visual contents of each cluster.
@@
To assess the model’s focus on specific neural networks, we utilized Grad-CAM [Selvaraju et al., 2020], a gradient-based technique. This method involves tracking the gradients from the target output back through the network layers to the originating fMRI input data. This process helps in pinpointing which areas of the input are most influential in the model’s predictions, where the gradient values are high. In our cluster analysis, the target outputs are the clusters. We created gradient-based saliency map for each cluster. As shown in Fig. 3, we averaged the gradient values from four subjects in the dataset, and we took the top 20 regions for visualization.
@@
During the training and reconstruction phases, we incorporated a visual region mask into the fMRI input data. This mask, aligned with the DiFuMo atlas [Dadi et al., 2020], sets the values within the visual regions to zero by following the network mapping scheme to the Yeo 7 networks Yeo et al. [2011]. Specifically, we applied the mask to all input fMRI data associated with the visual network as defined in the Yeo 7 classification. This masking procedure was consistently maintained throughout the training and image reconstruction processes. The images reconstructed from the fMRI data with the visual networks masked are displayed in Fig. 2. The ablation network ablation of other six networks are the same process described as visual network above.
@@ 4.4 Zero-Shot Human Imagination Decoding
To assess the quality of the reconstructed images against the ground truth texts, we utilized the Two-Way-Identification Accuracy metric. This involved using a pre-trained CLIP model to encode $n$ reconstructed images and $n$ ground-truth texts into image embeddings $\mathcal{H_{\textit{image}}}$ and text embeddings $\mathcal{H_{\textit{text}}}$, respectively. We computed the Pearson product-moment correlation coefficients Harris et al. [2020], resulting in an image-text coefficient matrix $M\in\mathbb{R}^{n\times n}$. The accuracy for each image was determined by the ratio of row elements with values less than their diagonal counterparts, divided by n1n-1. The overall two-way identification accuracy is the average of these individual accuracies.
We recruited 24 young adults (age under 30), each participant was presented with 20 scenarios accompanied by text descriptions. Utilizing the aforementioned metric, we aggregated the data into a comprehensive coefficient matrix $M_{\textit{all}}\in\mathbb{R}^{480\times 480}$, and performed a permutation test 5,000 times by randomizing the sequence of image embeddings.
@@
Acknowledgment
This research is supported by the HAI-Google Research Award, UST, HAI Hoffman-Yee Award, and National Institutes of Health grants NIH U24AG072701 and R01AG089169. We also greatly thank Yurong Liu from New York University for providing figure suggestions and proofreading.
Code and data availability
@@ A.1 Individual Decoding Result
In our main paper, we present two experiment settings, the individual model for decoding and the universal model. To explore our WAVE model decoding ability for individual subjects. We provide each subject metric results here. As shown in the Table 3. We showed all four subjects in the BOLD5000 dataset, and each subject ID is listed in the first column, and the details metrics results are followed in each row. WAVE has great advances in decoding semantic meaning from the fMRI and the results in High-Level and Semantic Top-1 Accuracy also prove this point. these results demonstrate that our model maintains high performance stability across all subjects (CSI1–CSI4). The inter-subject variance is notably low, particularly in high-level semantic metrics (e.g., CLIP and Top-1 Accuracy), where the model consistently outperforms baselines.
@@
[…]
Decoding from Non-Visual Networks.
Besides presenting the decoding results from whole brain wise data (WAVE) and visual cortex data (MindEye & Mind-Vis), we also explored decoding information from brain regions that do not belong to the visual networks. By applying the mask to the input fMRI data, WAVE can still successfully decode the natural image from fMRI across all subjects. In addition to the averaged subject results in the manuscript, we also provide the individual visual networks mask results in Table 4. Our individual visual mask result gives new insights into neuroscience to understand various brain regions’ functions not only at visual perception level, but in the cognition. These findings provide critical neuroscientific insights, suggesting that visual experience is not confined to perceptual areas but is deeply integrated into higher-order cognitive functions. As expected, the quantitative metrics for this masked condition were slightly lower than the unmasked whole-brain results reported in Table 3.
@@
[…]
@@
We conducted ablation studies to evaluate the performance of different encoder and decoder configurations in diffusion-based image reconstruction. Two types of encoders were used: the pre-trained fMRI Foundation model and the train-from-scratch MLP Backbone. For the decoders, we explored two diffusion models: Versatile Diffusion and Stable Diffusion. The ablation results, presented in Table 5, indicate that the fMRI Model, which employs a pretrained whole-brain fMRI encoder Thomas et al. [2022], demonstrates superior performance compared to the MLP Backbone. Additionally, Versatile Diffusion consistently outperformed Stable Diffusion across evaluated metrics.
@@
fMRI Encoder Decoder Acc(%) Min(%) Max(%) FID Min Max Region Enc Voxel Enc fMRI Model (Thomas et al. [2022]) w/o VD 22.13 ± 1.48 19.62 23.41 2.45 ± 0.71 1.68 3.55 fMRI Model (Thomas et al. [2022]) w/o SD 18.81 ± 1.17 17.87 20.46 3.70 ± 1.48 2.37 5.77 MLP Backbone w/o VD 18.79 ± 1.72 16.71 20.92 4.81 ± 0.95 3.53 5.80 MLP Backbone w/o SD 13.03 ± 0.95 12.17 14.35 6.69 ± 1.55 4.94 8.70 w/o MLP Backbone VD 12.91 ± 1.12 11.41 14.09 3.65 ± 0.63 3.09 4.53 w/o MLP Backbone SD 10.41 ± 1.00 9.01 11.23 6.62 ± 1.36 4.87 8.20
In addition to exploring various encoder and decoder configurations, our research extended to analyzing the impact of decoder loss functions on Subject CSI3. Our findings indicate that utilizing diffusion model loss directly on image pixel values, rather than on embeddings, enhances performance in low-level pixel-related metrics and Fréchet Inception Distance (FID). As demonstrated in Table 6, Mind-Vis Chen et al. [2023] has better low-level performance.
@@
Model Loss Decoder Acc(%) Min(%) Max(%) FID Min Max WAVE Embed VD 22.13 ± 1.48 19.62 23.41 2.45 ± 0.71 1.68 3.55 MindEye (Scotti et al. [2024]) Embed VD 13.45 ± 0.13 12.89 14.97 2.91 ± 0.39 2.53 3.30 Mind-Vis (Chen et al. [2023]) Pixel LDM 14.83 ± 2.25 12.59 19.08 1.64 ± 0.19 1.40 1.98
We also conduct research to show multi-modal fMRI-Visual-Language learning with prompt Meta-Net effect in this decoding task. As illustrated in Table 7, incorporating multiple modalities through contrastive learning enhances the model’s ability to encode representations, thereby improving the classification accuracy for Brain (fMRI) and Vision (Visual Stimulus) using the CLIP model. The introduction of prompt-based learning further aids in balancing the training across multiple modalities, enhancing both Brain-Vision and Brain-Language accuracy.
@@
Recent work has used linear methods like ridge regression for fMRI encoding Takagi and Nishimoto [2023]. To evaluate our pretrained fMRI encoder, we conduct two ablation studies. First, we test fMRI-to-image retrieval using CLIP cosine similarity between projected fMRI and image embeddings (Table 8). Second, we train encoders end-to-end on subject classification across all 4 subjects in BOLD5000 (Table 9). Both tasks demonstrate that (1) non-linear encoders substantially outperform linear baselines, and (2) pretraining provides significant additional gains, capturing both visual semantics and cross-subject variance.
Table 8: fMRI Encoder Ablation: Retrieval Accuracy. We compare three fMRI encoder configurations on Brain-Vision and Brain-Language retrieval tasks. The pretrained WAVE encoder outperforms both the non-pretrained version and linear baseline, demonstrating that autoregressive pretraining learns robust fMRI representations for semantic alignment.
Model Brain-Vision Accuracy (%) Brain-Language Accuracy (%) pretrained (WAVE) 53.98 36.28 w/o Pretrained 50.22 28.84 MLP backbone 47.36 26.57 Linear 40.78 25.67
Table 9: fMRI Encoder Ablation: Subject Classification Accuracy. We evaluate each encoder’s ability to capture cross-subject variance through subject classification on BOLD5000. The pretrained WAVE encoder achieves highest accuracy, showing that pretraining effectively learns cross-subject patterns in fMRI data.
Model Subject Classification (%) pretrained (WAVE) 84.78 w/o Pretrained 76.24 MLP backbone 74.85 Linear 68.32